### v1

@@ 5.3.1 Visual Interpretation using Grad-CAM
Grad-CAM [10] was applied to the layer preceding the softmax activation used for generating the predictions. This layer was particularly selected because it directly contributes to the decision-making process, offering a transparent view into how the model weighs different regions in its segmentation task. The visualizations from Grad-CAM, as shown in Figure 6, provide insights into the regions the models deemed most important in making their predictions. The images reveal that the attention mechanism seems to focus too intently on tumor regions, potentially overlooking broader contexts. Notably, while attention appears to work as designed by emphasizing certain regions, it might not always align with the most accurate predictions, especially with areas like the necrotic region being overemphasized. Additionally, the model sometimes misinterprets the texture of regions, resulting in potential misclassifications. This is particularly noticeable between necrotic and enhancing regions, which may have similar textures.
@@ 5.3.2 Confusion Matrix Analysis and Implications
To provide quantitative evidence supporting the findings from Grad-CAM visualizations, we perform a confusion matrix analysis, as shown in Figure 7. This matrix details how pixels from one class are misclassified as those from another, offering insights into the LATUP-Net model behavior. It is important to note that, there were patients with no enhanced tumor (ET) class in their ground truth. These cases have not been considered for the confusion matrix, as including them would distort the percentages. A primary observation is the misclassification between the necrotic and edema regions, a claim also derived from the qualitative visuals (see Section 5.3.1). The model perceives textural similarities between these areas, further complicated by nested tumor structures. Additionally, the necrotic region is notably vulnerable to misclassification, perhaps due to its texture and position within the tumor’s structure. When examining the standard deviation in Figure 8, combined with the average misclassification percentages, the classification performance variability is clear. This variability might spotlight outliers or deviations in model predictions, with classes like the necrotic region and enhancing region showing significant misclassification variability. The variations imply that certain samples heavily influence averages. In conclusion, the attention mechanism, while offering nuanced insights, can overly prioritize local features, causing the model to overlook important topological relationships recognized by human experts. A balance between local and global contexts seems important.
@@ 5.4 Comparison with State-of-the-Art Models
We compare our LATUP-Net model with the state-of-the-art models on the BraTS 2020 and 2021 datasets. We conduct five-fold cross-validation resulting in the performance results shown in Figures 9 and 10 of the test sets. We compare these results with other high-performing models and directly refer to related publications to obtain the evaluation results of the corresponding methods, which is a common approach in the study of brain tumor segmentation, as the source code of many existing brain tumor segmentation methods has not been released, and to avoid the bias introduced by model retraining.
@@ 5.4.1 Comparison with BraTS 2020 Results
We compare our LATUP-Net model with various state-of-the-art models on the BraTS 2020 dataset. The evaluation focuses on three key metrics: the Dice similarity coefficient (DSC), the percentile Hausdorff distance (HD95), and the number of model parameters. For our LATUP-Net model, we present the average performance across five-fold cross-validation to ensure a robust and comprehensive assessment. In contrast, for the state-of-the-art models, we report the results as they are presented in their respective original publications, which may include their best-split results or cross-validation outcomes. Detailed results of this comparative analysis are compiled in Table 4.
@@
Our model surpasses several others in DSC measurements for whole tumor segmentation. Specifically, it outperforms Raza et al. [43], Ballestar et al. [44], and Messaoudi et al. [45] by , , and respectively.
@@
Although a deeper analysis is needed to determine the exact mechanisms behind this efficiency, the results from the earlier model comparison section provide concrete evidence of the effectiveness of our approach. Specifically, our adoption of parallel convolutions in the first encoder block appears to play a crucial role. This is evidenced by the PC model having M fewer parameters than U-Net, yet achieving rapid convergence during the initial training epochs. This leads to shorter training times and a reduced need for computational resources.
@@
Brain tumor segmentation, particularly the tumor core and enhancing tumor, poses significant challenges. While models like nnU-Net exhibit strong results, our LATUP-Net model showcases similar scores. It is critical to differentiate between DSC and HD95 scores. Although our HD95 scores indicate accurate boundary predictions, our DSC for the enhancing tumor (ET) of is not the highest, signaling areas of improvement in volumetric overlap with the ground truth.
@@
Concerning the number of parameters, our model is remarkably efficient, with only M parameters, a stark contrast to the M parameters in nnU-Net. This efficiency reduces computational costs and the time required for processing, which is crucial for real-time clinical applications, while the performance remains comparable and even better in some aspects (see the HD95 results). Even if Although Raza et al. [43] has a parameter count closer to ours, they do not achieve a performance comparable to ours, highlighting the efficiency of our architecture.
@@ 5.4.2 Comparison with BraTS 2021 Results
Evaluating our LATUP-Net model on the more recent BraTS 2021 dataset shows promising results, comparable to state-of-the-art models (see Table 5). Specifically, our model achieves robust performance in segmenting the tumor core, with a DSC of . This surpasses many of the other models listed. However, it’s notable that Hatamizadeh et al. [52] displayed competitive DSC values, especially for WT and ET categories. Our model holds an advantage in terms of fewer parameters, making it computationally more efficient. The patterns observed in this dataset align closely with those from BraTS 2020, suggesting consistent performance.
@@
Our model’s DSC for the whole tumor and the enhanced tumor stand at and , respectively. When we mention these scores are in line with leading models, we specifically refer to Hatamizadeh et al. [52], Jia et al. [50], and Ma et al. [9], as shown in Table 5.
Considering HD95 values, our model consistently demonstrates accuracy in determining tumor boundaries. For the whole tumor, tumor core, and enhanced tumor, the HD95 scores are , , and , respectively. These figures, especially for the enhanced tumor, indicate that our model’s predictions closely align with actual tumor boundaries.
A significant advantage of the LATUP-Net model is its efficiency. With only M parameters, it is much lighter than Hatamizadeh et al.’s model [52] with M and Jia et al.’s model [50] with M. For models without a specified parameter count in the table, while we do not have the exact numbers, it is commonly understood in the field that such models typically have a considerable number of parameters due to their complex architectures.
In conclusion, the LATUP-Net model presents a fine balance between performance and efficiency, demonstrating the potential for practical clinical applications where both, accuracy and computational efficiency, are paramount.
6 Conclusion and Future Directions
@@
Our model demonstrates an impressive ability to delineate tumor boundaries with high accuracy, as evidenced by its performance in Hausdorff distance (HD95) measurements. These achievements indicate the model’s potential utility in clinical settings, where precise segmentations are integral to formulating effective treatment plans.
@@
The LATUP-Net model stands as a testament to the possibility of achieving state-of-the-art performance with a fraction of the computational cost, highlighting a promising direction for medical image analysis research and the development of practical, accessible tools for brain tumor segmentation. Future work includes adapting the architecture to other medical imaging segmentation tasks, and refining the balance between attention and convolutional features, especially to enhance our model’s sensitivity to enhancing tumor regions and to reduce the variations in segmentation performance across different regions.

### v2

@@ 5.4.1 Visual Interpretation using Grad-CAM
Grad-CAM [20] was applied to the layer preceding the softmax activation used for generating the predictions. This layer was particularly selected because it directly contributes to the decision-making process, offering a transparent view into how the model weighs different regions in its segmentation task. Grad-CAM visualizations, as shown in Figure 7, demonstrate that both models focus primarily on tumor areas. While the attention mechanism intensifies this focus, it does not always lead to the most accurate predictions. In particular, the necrotic region tends to be overemphasized, leading to misclassifications, as necrotic and enhancing regions often share similar textures.
This observation suggests that while attention appears to work as designed by emphasizing certain regions, it may not contribute to improved performance due to its reliance on local features. Grad-CAM further reveals that the segmentation task is heavily driven by these local features, like texture and boundaries, which attention does not necessarily enhance in a meaningful way. Additionally, because Grad-CAM tends to focus on the regions directly involved in segmentation, it does not provide insights into global relationships, such as left-right symmetry of the brain, which the attention mechanism might have been expected to capture.
Overall, the results indicate that the attention mechanism may have limited impact on segmentation performance, as the task is largely dependent on local feature extraction, and the attention mechanism does not provide significant additional context in this particular use case.
@@ 5.4.2 Confusion Matrix Analysis and Implications
To provide quantitative evidence supporting the findings from Grad-CAM visualizations, we perform a confusion matrix analysis, as shown in Figure 8. This matrix details how pixels from one class are misclassified as those from another, offering insights into the LATUP-Net model behavior. It is important to note that there were patients with no enhanced tumor (ET) class in the ground truth. These cases have not been considered for the confusion matrix, as including them would distort the percentages. A primary observation is the misclassification between the necrotic and edema regions, a claim also derived from the qualitative visuals (see Section 5.4.1). The model perceives textural similarities between these areas, further complicated by nested tumor structures. Additionally, the necrotic region is notably vulnerable to misclassification, perhaps due to its texture and position within the tumor’s structure. When examining the standard deviation in Figure 9, combined with the average misclassification percentages, the classification performance variability is clear. This variability might spotlight outliers or deviations in model predictions, with classes like the necrotic region and enhancing region showing significant misclassification variability. The variations imply that certain samples heavily influence averages. In conclusion, the attention mechanism, while offering nuanced insights, can overly prioritize local features, causing the model to overlook important topological relationships recognized by human experts. A balance between local and global contexts seems important.
@@ 5.5 Comparison with State-of-the-Art Models
We perform five-fold cross-validation, with results shown in Figures 10 and 11. These figures display the per-sample performance of our LATUP-Net model, highlighting the distribution of DSC and HD95 metrics, as well as variability, including outliers. Unlike other studies that report only average metrics, we provide a detailed analysis to demonstrate the model’s consistency across test samples.
We then compare the average results of our five-fold cross-validation with other high-performing models, as shown in Tables 5.5.1 and 5.5.1. For these comparisons, we rely on the evaluation results reported in related publications, a common practice in brain tumor segmentation research. This approach is necessary because the source code for many existing methods is unavailable, and it helps avoid the bias that could be introduced by retraining models.
@@ 5.5.1 Comparison with BraTS 2020 Results
We compare our LATUP-Net model with various state-of-the-art models on the BraTS 2020 dataset. The evaluation focuses on three key metrics: the Dice similarity coefficient (DSC), the percentile Hausdorff distance (HD95), and the number of model parameters, alongside GFLOPs to reflect computational complexity. For our LATUP-Net model, we present the average performance across five-fold cross-validation to ensure a robust and comprehensive assessment. In contrast, for the state-of-the-art models, we report the results as they are presented in their respective original publications, which may include their best-split results or cross-validation outcomes. Detailed results of this comparative analysis are compiled in Table 5.5.1.
@@
Our model surpasses several others in DSC measurements for whole tumor segmentation. Specifically, it outperforms Raza et al. [53], Ballestar et al. [54], and Messaoudi et al. [55] by , , and respectively. Additionally, it achieves a notable improvement in efficiency, having only M parameters compared to [44], [23], and [53]. Moreover, while Wang et al. [23] model requires GFLOPs, and Raza et al. [53] use GFLOPs, our model operates with only GFLOPs, highlighting its computational efficiency which is a crucial factor in real-world clinical applications.
When comparing the latest study by Zhu et al. [10], which demonstrates high performance with a DSC of for the whole tumor, our model’s is still competitive. However, LATUP-Net surpasses Zhu’s model in HD95, especially in tumor core and enhancing tumor predictions, demonstrating that it excels in boundary accuracy—a critical aspect in segmentation tasks. Although Zhu’s model outperforms in DSC across tumor regions, the computational complexity and parameter count are not mentioned. However, by incorporating transformer into the U-Net architecture, it is a moderately sized model compared to LATUP-Net, considering its performance trade-offs.
It is important to note that LATUP-Net’s superior performance is not due to common confounding factors, such as ensembling or leveraging additional data. This aligns with the recent recommendations [9], which cautions against practices that artificially inflate a model’s performance. Our architecture’s innovations—parallel convolution strategies and lightweight attention mechanisms—demonstrate genuine advancements in segmentation tasks without introducing unnecessary complexity or relying on inflated hardware resources. These results underscore the efficacy of simpler, well-configured models in achieving competitive performance in resource-constrained settings.
@@
Although a deeper analysis is needed to fully determine the mechanisms behind this efficiency, the results from the earlier model comparison section provide concrete evidence of the effectiveness of our approach. Specifically, our adoption of parallel convolutions in the first encoder block seems to play a crucial role. This is evident as the PC model has million fewer parameters than U-Net, yet achieves rapid convergence during the initial training epochs, leading to shorter training times and reduced computational needs.
@@
Brain tumor segmentation, particularly in the tumor core and enhancing tumor regions, poses significant challenges. While models like nnU-Net exhibit strong results, our LATUP-Net model achieves similar scores. It is critical to differentiate between DSC and HD95 scores. Although our HD95 scores indicate highly accurate boundary predictions, our DSC for the enhancing tumor (ET) is , which is not the highest, highlighting an area for improvement in volumetric overlap with the ground truth.
@@
In terms of the number of parameters, our model is remarkably efficient, with only million parameters, a stark contrast to nnU-Net’s million parameters. This efficiency reduces both computational costs and processing time, which is critical for real-time clinical applications, while maintaining comparable or even better performance in some aspects, as seen in the HD95 results. Although Raza et al. [53] has a parameter count closer to ours, their performance does not match ours, underscoring the effectiveness of our architecture.
rowsep=3pt
{tblr}
colspec=Q[2.5cm,valign=m]llllllllll,
cell1-21-9=gray!20,
cell11=r=2l,
cell12=c=2l,
cell14=c=2l,
cell16=c=2l,
cell18=r=2l,
cell19=r=2l
Study WT TC ET Parameters GFLOPs
DSC (%) HD95 (mm) DSC (%) HD95 (mm) DSC (%) HD95 (mm)
3D U-Net Baseline M
Isensee et al. [44] M -
Tang et al. [56] - -
Ballestar et al. [54] - -
Wang et al. [23] M 333
Messaoudi et al. [55] - - - - -
Raza et al. [53] - - - M 374.04
Zhu et al. [10] - -
LATUP-Net (proposed) M
rowsep=3pt
{tblr}
colspec=Q[2.5cm,valign=m]llllllllll,
cell1-21-9=gray!20,
cell11=r=2l,
cell12=c=2l,
cell14=c=2l,
cell16=c=2l,
cell18=r=2l,
cell19=r=2l
Study WT TC ET Parameters GFLOPs
DSC (%) HD95 (mm) DSC (%) HD95 (mm) DSC (%) HD95 (mm)
Peiris et al. [57] - -
Akbar et al. [58] - -
Jia et al. [59] M 449.79
Li et al. [60] - -
Ma et al. [61] - -
Hatamizadeh et al. [62] M
Roth et al. [63] - -
Zhu et al. [10] - -
LATUP-Net (proposed) M
@@ 5.5.2 Comparison with BraTS 2021 Results
Upon evaluating our LATUP-Net model on the BraTS 2021 dataset, it continues to show promising results, particularly in segmenting the tumor core, with a DSC of , outperforming several models. Compared to Hatamizadeh et al. [62], who also achieved competitive results across tumor regions, LATUP-Net demonstrates a clear advantage in computational efficiency. Our model has only million parameters and requires GFLOPs, significantly less than Hatamizadeh’s model, which requires million parameters and GFLOPs. This efficiency underscores the suitability of LATUP-Net for practical clinical use, especially in environments where computational resources may be limited.
In terms of HD95, LATUP-Net consistently achieves better results, with scores of for the whole tumor and for the tumor core. Zhu et al. [10], who achieved top DSC scores for the BraTS 2021 dataset, also produced competitive results in HD95. Nevertheless, LATUP-Net’s reduced parameter count and GFLOPs stand out as key differentiators, especially for real-time clinical applications where faster processing times are essential.
@@
Our model’s DSC for the whole tumor and the enhancing tumor are and , respectively. When we state that these scores are in line with leading models, we specifically refer to the work of Hatamizadeh et al., Jia et al. [59], and Ma et al. [61] (see Table 5.5.1).
In summary, the LATUP-Net model offers a fine balance between segmentation accuracy and computational efficiency. Despite its compact architecture, it delivers strong performance, particularly in HD95 metrics, which are crucial for precise tumor boundary delineation. With only million parameters and GFLOPs, LATUP-Net is highly suited for resource-constrained environments, offering a feasible solution for real-time clinical deployment. While models like Zhu et al. surpass LATUP-Net in certain DSC scores, the efficiency and competitive performance across HD95 metrics highlight the real-world applicability of our model in brain tumor segmentation tasks.
5.6 Limitations and Future Directions
Our model leverages multi-sequence MRI data for brain tumor segmentation and has demonstrated good performance on the BraTS 2020 and 2021 datasets, particularly when compared to other lightweight models working with similar data. However, we acknowledge that our work was trained on a relatively small dataset of about patients, which limits our understanding of the full covariance of potential data variations across different centers. This limitation is not unique to our study, as many segmentation models face similar challenges when dealing with multi-center data variations, especially regarding the availability of diverse data sources. We believe that access to larger, more heterogeneous datasets could help address these limitations, but this is an ongoing challenge within the field.
Other key limitations of our study lie in the consistency of model and parameter selection. While we have designed an effective architecture, the process of consistently selecting optimal models and fine-tuning hyperparameters has not been as rigorous as it could be. This challenge is partly due to the availability of computational resources, which restricts us from systematically testing a wider range of models and configurations. A more thorough exploration of model and parameter selection would be ideal, but this would require significant additional resources. As a result, we acknowledge that the model selection process could be improved and we plan to address this in future work.
Future work also includes adapting the architecture to other medical imaging segmentation tasks and refining the balance between attention and convolutional features, particularly to enhance our model’s sensitivity to tumor regions and reduce variations in segmentation performance across different regions. Another crucial direction will be to explore the model’s robustness when faced with incomplete or missing MRI modalities, a common scenario in real-world clinical settings. Ensuring that LATUP-Net can maintain high segmentation performance even when some modalities are unavailable will enhance its practical applicability in diverse medical environments. Additionally, we are currently studying the robustness and explainability of our model, which are critical for clinical applications [64]. However, this ongoing research extends beyond the scope of the present paper and will be addressed in future studies.
6 Conclusion
@@
Our model demonstrates an impressive ability to delineate tumor boundaries with high accuracy, as evidenced by its performance in Hausdorff distance (HD95) measurements. These achievements indicate the model’s potential utility in clinical settings, where precise segmentations are integral to formulating effective treatment plans. Furthermore, LATUP-Net’s lightweight architecture, requiring only GFLOPs, makes it particularly suitable for deployment in resource-constrained environments, such as developing countries, where computational resources may be limited.
@@
The LATUP-Net model stands as a testament to the possibility of achieving state-of-the-art performance with a fraction of the computational cost, highlighting a promising direction for medical image analysis research and the development of practical, accessible tools for brain tumor segmentation. In real-world clinical applications, however, the dependency on multi-sequence MRI data presents a practical challenge, as some modalities may be unavailable in certain settings. Addressing this limitation is a key consideration for future work.
Supplementary Material
7 Investigation of the Effect of Different Upsampling Approaches
In order to find the optimal upsampling method for the LATUP-Net architecture, we investigate two distinct techniques to identify the one that yields the best performance. Common upsampling approaches in lightweight networks are linear interpolation and transposed convolution, which are both tested here. Since our dataset consists of 3D MRI scans, trilinear interpolation is essential as it enables estimation across three dimensions by interpolating the values between neighboring voxels in a continuous manner. Trilinear interpolation preserves spatial relationships while performing upsampling and ensures that interpolated values smoothly transition between the feature mappings in the previous layer. Transposed convolution, instead, works by reversing the process of convolution to upsample the input, with the output size controlled by the kernel and stride parameters. Although transposed convolution can recover spatial resolution, it tends to introduce checkerboard artifacts, which can negatively impact segmentation performance. The experimental outcomes presented in Table S1 demonstrate that trilinear interpolation outperforms transposed convolution across all Dice scores, particularly in the tumor core (TC) and enhancing tumor (ET) regions. This demonstrates that, while both methods maintain same parameter efficiency, trilinear interpolation offers better precision in capturing finer details of tumor structures, making it the preferred upsampling strategy for LATUP-Net.
8 Comparison of Different Attention Mechanisms
We analyze the efficiency of various attention mechanisms in brain tumor segmentation on the BraTS 2020 test set. The evaluation is based on the Dice similarity coefficient (DSC) for whole tumor (WT), tumor core (TC), and enhancing tumor (ET) segmentation. Additionally, we provide the total training time for each model for reference, although note this information dependends strongly on the GPU hardware used (an RTX3060 in this case). Thus, model complexity comparisons based on GFLOPs, the number of parameters, and inference time are provided in Section 5.3, Table 5 of the main paper.
9 Investigation of the Effect of Different Loss Functions
In the LATUP-Net architecture, we evaluated multiple loss functions to determine the most effective one for optimizing the segmentation of 3D MRI data. The choice of loss function is critical as it directly influences the network’s ability to handle class imbalance and predict accurate segmentation. We experimented with several loss functions: Dice loss, Focal loss, Wasserstein Dice Loss, a combined Dice with Focal loss, weighted Dice loss with different weights for each class (WT, TC, ET), and weighted Dice loss with equal weights across all classes (WT, TC, ET).
The results summarized in Table S3 demonstrate that while each loss function has its strengths, the proposed weighted Dice loss (with class-specific weights) yields the best segmentation results.
10 Investigation of the Effect of Different Dropout Rates
Dropout is a widely used regularization technique that prevents overfitting by randomly dropping neurons during training, reducing reliance on any specific features and forcing the network to generalize better. We tested three dropout rates: , , and .
As shown in Table S4, a dropout rate of yielded the best performance, providing a balance between reducing overfitting and maintaining model capacity. The dropout rates of and performed slightly worse, likely due to too many neurons being dropped, leading to underfitting.
Table S1
Segmentation performance results for employing different upsampling strategies in LATUP-Net on the BraTS 2020 test set. WT – whole tumor, TC – tumor core, ET – enhancing tumor. Experiments were conducted using an RTX3060 GPU, after adding the proposed PC block, and the model was trained using the Dice loss function.
{tblr}
cell11 = r=2,
cell12 = r=2,
cell13 = c=3halign=c,
hline1,3-5 = -,
hline2 = 3-5,
Upsampling Strategy & Number of parameters Dice Score
WT TC ET
Convolution Transpose
Trilinear interpolation
Table S2
Efficiency analysis of various attention mechanisms for brain tumor segmentation on the BraTS 2020 test set. WT – whole tumor, TC – tumor core, ET – enhancing tumor. The table includes the Dice score and total training time for each model. Experiments were conducted using an RTX3060 GPU, after adding the proposed PC block, and the model was trained using the Dice loss function.
{tblr}
cell11 = r=2,
cell12 = c=3halign=c,
hline1,3 = -,
hline2 = 2-4,
Attention Mechanism & Dice Score
WT TC ET Total Training Time (hours)
PC+SE
PC+CBAM
PC+CBAM+SE
PC+(SE-3D)
PC+Residual SE
PC+ECA
Table S3
Segmentation performance results for employing different loss functions in LATUP-Net on the BraTS 2020 test set. WT – whole tumor, TC – tumor core, ET – enhancing tumor. Experiments were conducted using an RTX3060 GPU, after adding the proposed PC block.
{tblr}
cell11 = r=2,
cell12 = c=3halign=c,
hline1,3 = -,
hline2 = 2-4,
Loss Function Dice Score
WT TC ET
Dice Loss
Focal Loss
Wasserstein Dice Loss
Dice + Focal Loss
Weighted Dice Loss (Different Weights)
Weighted Dice Loss (Equal Weights)
The proposed WDL
11 Investigation of the Effect of Different Regularizers
Regularization is another important technique to prevent overfitting. We experimented with L1, L2, and a combination of both L1 and L2 (Elastic Net) regularization. L1 regularization encourages sparsity, while L2 regularization penalizes large weights. We tested different values for each type of regularization: L1 with , weights, and L2 with , weights, and L1 + L2 with weights .
As summarized in Table S5, L2 regularization with a weight of resulted in the best segmentation performance, providing significant benefits in terms of preventing overfitting, as observed in the loss curve. The combination of L1 and L2 regularization with values also performed well, while stronger L1 regularization with weight reduced performance slightly.
Table S4
Segmentation performance results for employing different dropout rates in LATUP-Net on the BraTS 2020 test set. WT – whole tumor, TC – tumor core, ET – enhancing tumor. Experiments were conducted using an RTX3060 GPU, after adding the proposed PC block, and the model was trained using the Dice loss function.
{tblr}
cell11 = r=2,
cell12 = c=3halign=c,
hline1,3 = -,
hline2 = 2-4,
Dropout Rate Dice Score
WT TC ET
Table S5
Segmentation performance results for employing different regularizers in LATUP-Net on the BraTS 2020 test set. WT – whole tumor, TC – tumor core, ET – enhancing tumor. Experiments were conducted using an RTX3060 GPU, after adding the proposed PC block, and the model was trained using the Dice loss function.
{tblr}
cell11 = r=2,
cell12 = c=3halign=c,
hline1,3 = -,
hline2 = 2-4,
Regularizer, weight Dice Score
WT TC ET
L1,
L1,
L2,
L2,
(L1,L2),